2000mg of salt is mixed with 50ml of water how much slat per ml of water
Answer by jim_thompson5910(35256)

(2000 mg of salt)/(50 ml of water) = 2000/50 = 40

There is 40 mg of salt for every ml of water

In other words, 1 ml of water contains 40 mg of salt (apply this to any/every 1 ml unit of water).